# Discovering universal temperature regulation dynamics in animals

### v2

@@ 3.5 Towards a multi-scale model of seasonal physiology
$\{\text{dimensionless body temperature}\}~~\dot{\tilde{x}}$ $=\omega\tilde{x}^{2}-\tilde{x}^{3}-\nu(\tilde{x}\tilde{y}-\tilde{x}^{2}\tilde{y})+\zeta,$ $\tilde{x}(0)=\tilde{x}_{0},$ (5)
$\{\text{dimensionless unknown physiological state}\}~~\dot{\tilde{y}}$ $=\rho(\tilde{x}^{2}-\tilde{y}^{2}),$ $\tilde{y}(0)=\tilde{y}_{0}.$ (6)
Each of the parameters has a clear physiological interpretation. The parameter $\omega$ controls the amplitude of the oscillations during hibernation. From laboratory experiments, it is known that hibernators attenuate the amplitude of the oscillations in response to environmental factors [37]. The parameter $\nu$ controls the seasonally varying coupling between body temperature and the unknown physiological state. $\nu$ is the parameter that controls the transitions from homeothermy to poikilothermy and vice versa. Biologically, $\nu$ can be interpreted as a summary output of the circannual clock (see below). The parameters $\zeta$ and $\rho$ influence the period of oscillations during hibernation. $\rho$ controls the scale separation between body temperature and the unknown physiological state and influences the period of oscillations. Biologically, we can interpret $\zeta$ and $\rho$ as modulating the speed of the hourglass and the location of the threshold, perhaps in an environmentally sensitive manner [37].
The system undergoes a supercritical Hopf bifurcation as a function of $\nu$, and we can find where this bifurcation occurs in $\omega$-$\nu$-$\zeta$-$\rho$ space.  [^4] It is straightforward to analytically find where this bifurcation occurs in $\omega$-$\nu$-$\zeta$-$\rho$ space. This curve divides parameter space into qualitatively different behaviors and has the form $\nu^{4}-8\nu^{3}\zeta+2\nu^{3}\rho-4\nu^{3}\omega-4\nu^{2}\rho^{2}+4\nu^{2}\omega^{2}-8\nu\rho^{3}+16\nu\rho^{2}\omega-8\nu\rho\omega^{2}-2\nu^{3}+$ (7) $36\nu^{2}\zeta-6\nu^{2}\rho+9\nu^{2}\omega+12\nu\rho\omega-12\nu\omega^{2}+8\rho^{3}-12\rho^{2}\omega+4\omega^{3}-54\nu\zeta+27\zeta=0.$ The curve is visualized in $\omega$-$\nu$ space for fixed values of $\zeta$ and $\rho$ (see Figure 4 A). The transition from a regulated body temperature to an oscillating body temperature during hibernation is associated with a change of stability in the system (see Section 5) and traversing a boundary in parameter space (see Figure 4). Similarly, the exit from hibernation and reversion to a regulated body temperature exhibited in the summer is associated with the opposite change of stability and movement in parameter space (see Section 5 and Figure 4A). Assuming $\nu$ is time-dependent, we can find an example trajectory of $\nu(t)$ that allows the system to capture the broad features of seasonal hibernation physiology (Figure 4 B and C).
@@ 3.6 Thermophysiological organization across species
Here, we show the possibility of a conserved thermophysiological mechanism by modifying our discovered core model (Equations (10)-(11)) to respond to environmental information. To determine how to add environmental information to our models, we first examine the dimensionality of the thermophysiological system across species using time delay embedding. We find that the dimensionality of the underlying thermophysiological systems for the elephant shrew, noisy miner, and black bear is relatively low-dimensional across species (all systems appear to be three-dimensional, see Figure 5C and Sections 6.2-6.3). To approximate the extra dimensionality, we add environmental information to our core model. Mathematically, these take the form of trigonometric forcing functions. We find that our modified models can qualitatively describe the body temperature patterns of “daily-heterotherm” species (the noisy miner), an ”intermediate” species (the elephant shrew), and a large ”hibernator” (the black bear) (Figure 5B and D).
@@
To capture the thermoregulatory behavior of the black bear, we also add environmental information through forcing functions, which, again, modify the phase plane dynamics and limit cycle shape in a temporal manner. In this case, we find that we must add two time-dependent forcing functions to our model (Figure 5E, rightmost). Notably, the period of these forcing functions is much longer than the forcing functions added to the noisy miner and elephant shrew models, which are assumed to be circadian. Biologically, it is plausible that these forcing functions correspond to an entity controlled by the circannual clock, such as an endocrine molecule with slow, seasonal fluctuations. It is also possible that an entity under circannual control could have an interaction with a circadian process, as circadian gene expression was observed in grizzly bears and may also occur in black bears [38].
Figure 5: We modify our discovered physiological model of Arctic ground squirrel to include environmental information and find that we can qualitatively match body temperature data recorded from a wide range of species, including a bird, an elephant shrew, and a bear, which also dynamically modulate body temperature. This suggests that a universal, environmentally sensitive mechanism could regulate body temperature across a diverse range of species. A: Photographs of an Arctic ground squirrel, an elephant shrew, a noisy miner, and a black bear. All images taken from Wikipedia. B: Temperature recordings from an Arctic ground squirrel [9], an elephant shrew [23], a noisy miner [39], and a black bear [21]. C: Phase space reconstructions of underlying thermophysiological systems. These systems are 2-to 3-dimensional. D: Model simulations of body temperature across species (blue) qualitatively match the observed body temperature data in B. Inferred dynamics of a hidden physiological state (pink). E: Environmental information interacting with the core model for each species.
We find we can qualitatively describe a vast range of body temperature patterns across species using a core thermophysiological model modified to respond to different types of environmental information. Small hibernators use just the core mechanism during hibernation with no environmental input, birds and small non-hibernating mammals use the core mechanism modified to include a circadian-dependent signal, and large hibernators use the core mechanism modified to include a circannual-dependent signal. This is a more mechanistic restructuring of our current biological understanding of thermophysiology across species. This analysis leads to a new hypothesis: a core thermophysiological mechanism is conserved across heterothermic endotherms, which can respond to environmental signals (Figure 5).
@@ 4 Discussion
Here, we have used a blend of automated model discovery and classical dynamical systems theory to differentiate between two physiological mechanisms hypothesized to drive interbout arousals during small-mammalian hibernation and comment on the possible organization of thermophysiological systems across species using heterothermic endothermy. We find evidence that interbout arousals are driven by an Hourglass and Threshold-type mechanism, as hypothesized in [8] and work towards a multi-scale understanding of seasonal physiological decision making through a bifurcation analysis (see Section 3.5). We modified our discovered core model to include environmental signals and found that we could qualitatively match body temperature recordings from a variety of species, including a bird, an elephant shrew, and a black bear (Figure 5). This leads to a new hypothesis: a core thermophysiological mechanism is conserved across heterothermic endotherms, which can respond to environmental signals.
@@
While the model we find is convincing in demonstrating that a core model could explain body temperature regulation across species, it is not unique. In addition to other structures that arise from symmetry transformations (see Section 3.4), there are likely other dynamical systems models that could adequately describe the thermophysiology of the Arctic ground squirrel. Some models that emerge from our model selection procedure do not immediately exhibit limit cycle behavior upon forward simulation. In these cases, it was difficult to know which terms need to be added “back” to the model to generate a limit cycle, or, in the case of poor fits (see Figure 2B, far off the Pareto front), if a limit cycle was even possible. Recall that there is no definitive technique for ruling out limit cycle behavior for general dynamical systems models, though it is possible in special cases (for examples of such systems see [40]). Mathematically, it would be of interest to determine if it is possible to rule out limit cycle behavior for any broader class of planar Ordinary Differential Equations and study additional model structures consistent with this data. As we cannot definitively determine the full class of models that can describe the thermophysiology of the Arctic ground squirrel for a single data set, we did not attempt to understand the possible variety of models that emerge if we examined other body temperature data sets from different Arctic ground squirrels. We successfully found and analyzed a data-driven model of Arctic ground squirrel thermophysiology, but future work will likely find alternative or modified model structures that can produce similar behavior.
@@
Beyond the algorithmic design, we made several practical modeling choices. We included monomials up to order three in our library during model selection. We drew inspiration from classical dynamical systems models from neuroscience, which typically include terms up to third degree, such as the FitzHugh-Nagumo Model [14, 15]. Models with higher-order terms, such as quartic and quintic terms, may work as well or better but would be more computationally challenging to select. Another choice was which term to add to the selected models to constrain the period of the limit cycle. We chose to add a small constant to the body temperature equation as the simplest option, but other small linear terms would also regulate the period length of the limit cycle. In future work, we wish to understand if our models can be transformed into Líenard systems or if we have discovered any candidates from a broader and distinct class of oscillatory models. In the qualitative modeling, we did not examine all possible ways to add environmental information. We demonstrate the existence of one version of a core model and a suite of extended, environmentally driven models, but we do not expect uniqueness. Similarly, it is important to note that the hidden state in our models of thermophysiology across species may not correspond to the same underlying biological entity in each species. We simply claim that the thermophysiological systems may have core components that interact in the same fashion across species, whether due to genetic conservation or convergence.
@@
Despite these limitations and caveats, our discovered macro-scale physiological model can be used to guide the ongoing search for molecular mechanisms that underpin small mammalian hibernation physiology based on the dynamics of the hidden state. However, our model cannot differentiate between an accumulating or depleting entity, or whether the driver is a single entity or an amalgamation of multiple entities. In practice, the causative agent may take the form of a single metabolite, the ratio of key metabolites, a difference between key metabolites, or something else entirely. Our model of hibernation physiology would be validated if the currently unknown causative agent oscillates with the same period as body temperature dynamics during hibernation and is constant during summer, when the organism exhibits a stable body temperature, or invalidated if not. A key point is that the causative agent may not be low in the summer. It may be relatively high, but we expect it to be constant. It is important to note that our qualitative modeling presented in see Section 3.5 captures only the broad features of the seasonal body temperature time series and was not intended to model the “flickering” dynamics that occur at the entrance and exit of hibernation [12]. In future work, we will modify the model to better capture these dynamics, but this was not the intention of this modeling effort.
@@
We discover a new data-driven physiological model in the style of the Fitzhugh-Nagumo model [14, 15], the NPZ model [16], and the Brusselator model [17, 18, 19], which distill the incredible complexity and retain only the essential features of the underlying system. We used our discovered model to provide new insights into hibernation physiology, posit the beginnings of a multi-scale model of hibernation that spans the neurological and physiological scales, and put forward a hypothesis in thermal biology: a core thermophysiological mechanism is conserved across heterothermic endotherms, which can respond to environmental signals. Our analysis starts to describe how physiological systems are affected and controlled by the external environment, and as more physiology is uncovered, we can further refine our models. In the future, we hope to further understand which features of the external environment, such as photoperiod, ambient temperature (see Figure 2 in [41]), magnetic fields, social interactions, and food availability, influence thermophysiology and physiology more broadly in potentially nonlinear and seasonally varying ways.
5 Appendix
The Arctic ground squirrel thermophysiology model we discovered (Eqs. 3-4) can also describe seasonal transitions between summer homeothermy and winter poikilothermy. In the summer, the Arctic ground squirrel is homeothermic, with minor fluctuations induced by activity and sleep around a physiological fixed point near 37 C. During hibernation, the organism cannot remain solely at 37 C or the low state (near or below 0 $\circ$ C), but must oscillate between the two states to survive. Therefore, any physiological fixed points corresponding to high-temperature or low-temperature states must be unstable (or non-attractive). Mathematically, we expect the stability of the model’s high-temperature fixed point to transition from stable to unstable in the fall and back to stable in the spring. The model’s low-temperature fixed point should always be unstable, as staying at a low temperature without oscillations at any time of year would be fatal for the organism. To confirm this, we perform a stability analysis of our model.
$\{\text{dimensionless body temperature}\}~~\dot{\hat{x}}$ $=\hat{x}^{2}-\hat{x}^{3}-\gamma\hat{x}\hat{y}+\xi\hat{x}^{2}\hat{y}+\iota,\hat{x}(0)=\hat{x}_{0},$ (8)
$\{\text{dimensionless unknown physiological state}\}~~\dot{\hat{y}}$ $=\kappa\left(\hat{x}^{2}-\hat{y}^{2}\right),\hat{y}(0)=\hat{y}_{0}.$ (9)
$\{\text{dimensionless body temperature}\}~~\dot{\hat{x}}$ $=\hat{x}\left(\hat{x}-\gamma\hat{y}\right)\left(1-\hat{x}\right)+\iota,\hat{x}(0)=\hat{x}_{0},$ (10)
$\{\text{dimensionless unknown physiological state}\}~~\dot{\hat{y}}$ $=\kappa\left(\hat{x}^{2}-\hat{y}^{2}\right),\hat{y}(0)=\hat{y}_{0}.$ (11)
The model parameters of Eqs. (10)-(11) has interpretable physiological meaning. $\gamma$ describes the seasonal coupling between body temperature and the unknown physiological state. The parameters $\iota$ and $\kappa$ influence the period of oscillations during hibernation. Eqs. (10)-(11) has three fixed points: $(\hat{x},\hat{y})=(0,0)$, (1,1)(1,-1), and $(1,1)$ when $\iota=0.$ [^5] It is reasonable to take $\iota=0$ to examine fixed points sufficiently far away from the origin, as the higher order terms will locally dominate $\iota$, as $\iota<<1$, but this is not reasonable near the origin where $\iota$ dominates. So under the $\iota=0$ approximation, we cannot comment on the stability of the fixed point at the origin, as our assumption does not hold. Additionally, note that the fixed point $(\hat{x},\hat{y})=(1,-1)$ is unstable for all $\gamma$ and $\kappa$. A linear stability analysis of Eqs. (10)-(11) when $\iota=0$ shows that $(\hat{x},\hat{y})=(1,1)$, which corresponds to the high-temperature state, is stable for $0<\gamma<1$ and is an unstable saddle for $\gamma>1$. For $\gamma=1$, linear stability analysis is indeterminate. Biologically, the thermoregulatory transitions between homeothermy and poikilothermy can be understood as seasonal variation in $\gamma$, perhaps driven by the circannual clock, which corresponds mathematically to a seasonal change of stability of the fixed point $(\hat{x},\hat{y})=(1,1)$, the high-temperature fixed point. While the Arctic ground squirrel is homeothermic during the summer ($\gamma<1$), the thermophysiological system stays near $(\hat{x},\hat{y})=(1,1)$ with minor fluctuations. During hibernation ($\gamma>1$), $(\hat{x},\hat{y})=(1,1)$ is unstable, and the system snaps to the limit cycle, which describes the poikilothermic regime. Thus, the stability of the point near (1,1) is controlled by $\gamma$ in a seasonal manner as expected. This seasonal transition and change of stability are shown in Figure 4 for the same model with a different choice of nondimensionalization.
$1<\gamma<2(1+\kappa)$ (12)

### v4

@@ 2.5 Towards a multi-scale model of seasonal physiology
$\{\text{dimensionless body temperature}\}\penalty 10000\ \penalty 10000\ \dot{\tilde{x}}$ $=\omega\tilde{x}^{2}-\tilde{x}^{3}-\nu(\tilde{x}\tilde{y}-\tilde{x}^{2}\tilde{y})+\zeta,$ $\tilde{x}(0)=\tilde{x}_{0},$ (5)
$\{\text{dimensionless unknown physiological state}\}\penalty 10000\ \penalty 10000\ \dot{\tilde{y}}$ $=\rho(\tilde{x}^{2}-\tilde{y}^{2}),$ $\tilde{y}(0)=\tilde{y}_{0}.$ (6)
Each of the parameters has a clear physiological interpretation. The parameter $\omega$ controls the amplitude of the oscillations during hibernation. From laboratory experiments, it is known that hibernators attenuate the amplitude of the oscillations in response to environmental factors [41]. It was also shown that perturbing temperature during hibernation can alter the timing of reproduction expression but does not interfere with the timing of future hibernation cycles [42]. The parameter $\nu$ controls the seasonally varying coupling between body temperature and the unknown physiological state. $\nu$ is the parameter that controls the transitions from homeothermy to the hibernation thermoregulatory regime and vice versa. Biologically, $\nu$ can be interpreted as a summary output of the circannual clock (see below). The parameters $\zeta$ and $\rho$ influence the period of oscillations during hibernation. $\rho$ controls the scale separation between body temperature and the unknown physiological state and influences the period of oscillations. Biologically, we can interpret $\zeta$ and $\rho$ as modulating the speed of the hourglass and the location of the threshold, perhaps in an environmentally sensitive manner [41].
The system undergoes a supercritical Hopf bifurcation as a function of $\nu$, and we can find where this bifurcation occurs in $\omega$-$\nu$-$\zeta$-$\rho$ space.  [^5] It is straightforward to analytically find where this bifurcation occurs in $\omega$-$\nu$-$\zeta$-$\rho$ space. This curve divides parameter space into qualitatively different behaviors and has the form $\nu^{4}-8\nu^{3}\zeta+2\nu^{3}\rho-4\nu^{3}\omega-4\nu^{2}\rho^{2}+4\nu^{2}\omega^{2}-8\nu\rho^{3}+16\nu\rho^{2}\omega-8\nu\rho\omega^{2}-2\nu^{3}+$ (7) $36\nu^{2}\zeta-6\nu^{2}\rho+9\nu^{2}\omega+12\nu\rho\omega-12\nu\omega^{2}+8\rho^{3}-12\rho^{2}\omega+4\omega^{3}-54\nu\zeta+27\zeta=0.$ The curve is visualized in $\omega$-$\nu$ space for fixed values of $\zeta$ and $\rho$ (see Figure 4 A). The transition from a regulated body temperature to an oscillating body temperature during hibernation is associated with a change of stability in the system (see Section 5.1) and traversing a boundary in parameter space (see Figure 4). Similarly, the exit from hibernation and reversion to a regulated body temperature exhibited in the summer is associated with the opposite change of stability and movement in parameter space (see Section 5.1 and Figure 4A). Assuming $\nu$ is time-dependent, we can find an example trajectory of $\nu(t)$ that allows the system to capture the broad features of seasonal hibernation physiology (Figure 4 B and C).
@@ 2.6 Thermophysiological organization across species
Here, we show the possibility of a conserved thermophysiological mechanism by modifying our discovered core model (Equations (10)-(11)) to respond to environmental information. To determine how to add environmental information to our models, we first examine the dimensionality of the thermophysiological system across species using time delay embedding. We find that the dimensionality of the underlying thermophysiological systems for the elephant shrew, noisy miner, and black bear is relatively low-dimensional across species (all systems appear to be three-dimensional, see Figure 5C and Sections 4.2-4.3). To approximate the extra dimensionality, we add environmental information to our core model. Mathematically, these take the form of trigonometric forcing functions. These functions are phenomenological placeholders, not inferred physiological processes. We find that our modified models can qualitatively describe the body temperature patterns of “daily-heterotherm” species (the noisy miner), an ”intermediate” species (the elephant shrew), and a large ”hibernator” (the black bear) (Figure 5B and D).
@@
To capture the thermoregulatory behavior of the black bear, we also add environmental information through forcing functions, which, again, modify the phase plane dynamics and limit cycle shape in a temporal manner. In this case, we find that we must add two time-dependent forcing functions to our model (Figure 5E, rightmost). Notably, the period of these forcing functions is much longer than the forcing functions added to the noisy miner and elephant shrew models, which are assumed to be circadian. Biologically, it is plausible that these forcing functions correspond to an entity oscillating on year scale, such as an endocrine molecule with slow, seasonal fluctuations. It is also possible that such an entity could interact with a circadian process, as circadian gene expression was observed in grizzly bears and may also occur in black bears [43].
Figure 5: We modify our discovered physiological model of Arctic ground squirrel to include environmental information and find that we can qualitatively match body temperature data recorded from a wide range of species, including a bird, an elephant shrew, and a bear, which also dynamically modulate body temperature. This suggests that a low-dimensional, environmentally sensitive mechanism could regulate body temperature across a diverse range of species. A: Photographs of an Arctic ground squirrel, an elephant shrew, a noisy miner, and a black bear. All images taken from Wikipedia. B: Temperature recordings from an Arctic ground squirrel [9], an elephant shrew [23], a noisy miner [44], and a black bear [21]. C: Phase space reconstructions of underlying thermophysiological systems. These systems are 2-to 3-dimensional. D: Model simulations of body temperature across species (blue) qualitatively match the observed body temperature data in B. Inferred dynamics of a hidden physiological state (pink). E: Internally-encoded environmental information interacting with the core model for each species. Time is dimensionless in panels D and E.
We find we can qualitatively describe a vast range of body temperature patterns across species using a core thermophysiological model modified to respond to different types of environmental information. Small hibernators use just the core mechanism during hibernation with no environmental input, birds and small non-hibernating mammals use the core mechanism modified to include a circadian-dependent signal, and large hibernators use the core mechanism modified to include signals that oscillate with a period much longer than that of the circadian system. Our analysis leads to a more mechanistic understanding of thermophysiology across species, and a new hypothesis: a core thermophysiological mechanism is shared across heterothermic endotherms, which can respond to environmental signals (Figures 5-6).
@@ 3 Discussion
Here, we have used a blend of automated model discovery and classical dynamical systems theory to differentiate between two physiological mechanisms hypothesized to drive interbout arousals during small-mammalian hibernation and comment on the possible organization of thermophysiological systems across species using heterothermic endothermy. We find evidence that interbout arousals are driven by hourglass-like dynamics, as hypothesized in [8] and work towards a multi-scale understanding of seasonal physiological decision making through a bifurcation analysis (see Section 2.5). We modified our discovered core model to include environmental signals and found that we could qualitatively match body temperature recordings from a variety of species, including a bird, an elephant shrew, and a black bear (Figure 5). This leads to a new hypothesis: a core thermophysiological mechanism is shared across heterothermic endotherms, which can respond to environmental signals (see Figure 6).
Figure 6: We hypothesize that a low-dimensional, environmentally sensitive core regulator could control body temperature across a diverse range of species. A broad range of external inputs including photoperiod, ambient temperature, social behavior, food availability, seasonal change, and other cues are neurally-integrated and represented as internal signals, such as the circannual output, a short period or circadian oscillation, or a long period oscillation. The internal signals interact with the core temperature control architecture and produce varied physiological output observed across different species.
@@
While the model we find is convincing in demonstrating that a core model could explain body temperature regulation across species, it is not unique. In addition to other structures that arise from symmetry transformations (see Section 2.4), there are likely other dynamical systems models that could adequately describe the thermophysiology of the Arctic ground squirrel. Some models that emerge from our model selection procedure do not immediately exhibit limit cycle behavior upon forward simulation. In these cases, it was difficult to know which terms need to be added “back” to the model to generate a limit cycle, or, in the case of poor fits (see Figure 2B, far off the Pareto front), if a limit cycle was even possible. Recall that there is no definitive technique for ruling out limit cycle behavior for general dynamical systems models, though it is possible in special cases (for examples of such systems see [38]). Mathematically, it would be of interest to determine if it is possible to rule out limit cycle behavior for any broader class of planar Ordinary Differential Equations and study additional model structures consistent with this data. To limit the computational burden of the initial model selection and given the fact that we cannot determine determine the full class of models that can describe the thermophysiology of the Arctic ground squirrel for a single data set, we did not attempt to understand the possible variety of models that emerge if we examined other body temperature data sets from different Arctic ground squirrels. We successfully found and analyzed a data-driven model of Arctic ground squirrel thermophysiology, but future work will likely find alternative or modified model structures that can produce similar behavior. However, we do expect our discovered model to generalize well across data from distinct individuals. Body temperature dynamics across individual squirrels are remarkably similar (see Figure 1 in [45]), and the existing subtle variations could be easily captured by different parameterizations of Eqs. (5)-(6).
@@
Beyond the algorithmic design, we made several practical modeling choices that should be noted. We included monomials up to order three in our library during model selection. We drew inspiration from classical dynamical systems models from neuroscience, which typically include terms up to third degree, such as the FitzHugh-Nagumo Model [14, 15]. Models with higher-order terms, such as quartic and quintic terms, may work as well or better but would be more computationally challenging to select. Another choice was which term to add to the selected models to constrain the period of the limit cycle. We chose to add a small constant to the body temperature equation as the simplest option, but other small linear terms would also regulate the period length of the limit cycle. In future work, we wish to understand if our models can be transformed into Líenard systems or if we have discovered any candidates from a broader and distinct class of oscillatory models. While we evaluated the fit of our models through multiple lenses, there is always the risk of overfitting the parameters, given the limited oscillatory data we used for training and validation. Given this challenge, it is surprising that we were able to discover a model that captures the oscillatory Arctic ground squirrel body temperature data (Section 2.3), capture broad seasonal transitions between thermoregulatory modes (Section 2.5), and generalizes across multiple different species (Section 2.6). In the qualitative modeling, we did not examine all possible ways to add environmental information. We demonstrate the existence of one version of a core model and a suite of extended, environmentally driven models, but we do not expect uniqueness. Similarly, it is important to note that the hidden state in our models of thermophysiology across species may not correspond to the same underlying biological entity in each species. We simply claim that the thermophysiological systems may have core components that interact in the same fashion across species, whether due to genetic conservation or convergence, both of which are plausible.
@@
Despite these limitations and caveats, our discovered macro-scale physiological model can be used to guide the ongoing search for molecular mechanisms that underpin small mammalian hibernation physiology based on the dynamics of the hidden state. Due to a scaling symmetry in our model, we cannot differentiate between an accumulating or depleting entity, or whether the driver is a single entity or an amalgamation of multiple entities (see Section 2.4). In practice, the causative agent may take the form of a single metabolite, the ratio of key metabolites, the difference between key metabolites, or something else entirely. Our model of hibernation physiology would be validated if the currently unknown causative agent oscillates with the same period as body temperature dynamics during hibernation and is constant during summer, when the organism exhibits a stable body temperature, or invalidated if not. A key point is that the causative agent may not be low in the summer. It may be relatively high, but we expect it to be constant. Functionally, experimental tests of our model would take the form of metabolomic or endocrine sampling over the course of the annual cycle and a demonstration of the causative nature of the discovered entity on interbout arousals. It is important to note that our qualitative modeling presented in see Section 2.5 captures only the broad features of the seasonal body temperature time series and was not intended to model the “flickering” dynamics that occur at the entrance and exit of hibernation [12]. In future work, we will modify the model to better capture these dynamics, but this was not the intention of this modeling effort.
@@
We discover a new data-driven physiological model in the style of the Fitzhugh-Nagumo model [14, 15], the NPZ model [16], and the Brusselator model [17, 18, 19], which distill the incredible complexity and retain only the essential features of the underlying system. We used our discovered model to provide new insights into hibernation physiology, posit the beginnings of a multi-scale model of hibernation that spans the neurological and physiological scales, and put forward a hypothesis in thermal biology: a core thermophysiological mechanism is shared across heterothermic endotherms, which can respond to environmental signals (see Figure 6). Our analysis starts to describe how physiological systems are affected and controlled by the external environment, and as more physiology is uncovered, we can further refine our models. In the future, we hope to further understand which features of the external environment, such as photoperiod, ambient temperature (see Figure 2 in [46]), magnetic fields, social interactions, and food availability, influence thermophysiology and physiology more broadly in potentially nonlinear and seasonally varying ways.